ABC Company employs a periodic inventory system and sells its inventory to customers
for $27 per unit. ABC Company had the following inventory information available for
the month of May:

May 1    Beginning inventory 2,400 units @ $14 cost per unit
May 8    Sold 1,000 units
May 13   Purchased 1,300 units @ $17 cost per unit
May 18   Sold 1,200 units
May 21   Purchased 1,100 units @ $11 cost per unit
May 28   Sold 300 units
May 30   Purchased 700 units @ $10 cost per unit

During May, ABC Company reported operating expenses of $14,500 and had an income tax
rate of 43%.

Calculate the dollar amount of ending inventory shown on ABC's May 31 balance sheet
using the FIFO method.
Answer #1

According to FIFO method, ending inventory consists of the units produced at the latest.

Number of units sold = 1,000+1,200+300

= 2,500 units

Number of units available for sale = Beginning inventory + Inventory purchases

= 2,400+(1,300+1,100+700+

= 5,500 units

Ending inventory units = Number of units available for sale - Number of units sold

= 5,500-2,500

= 3,000 units

Calculation of Ending Inventory Cost
Date Units Unit Cost Total Unit
May 13 1,200 17 20,400
May 21 1,100 11 12,100
May 30 700 10 7,000
3,000 $39,500

Ending inventory = $39,500

Kindly comment if you need further assistance.

Thanks‼!